# Complexity of Linear Subsequences of $k$-Automatic Sequences

### 07 Mar 2026 v3

@@ 5 Construction by Büchi Arithmetic
We use the total number of transitions of an automaton $M$ as a proxy for the computational complexity of constructing $M$. Sometimes the complexity bounds in this section include the addition of $O(\log i)$ or $O(i^{j})$ components for some constant $i$; these addition components are not significant compared to multiplication in $O(\log i)$ or $O(i^{j})$ components for some constant $i$.
@@
Furthermore, the minimization steps in the implementation for an $n$-state automaton can use an algorithm running in $O(n\log n)$ time by Hopcroft [16] or Valmari [26]. The algorithm by Hopcroft can be slightly modified to be applied to DFAOs as described by van Spaendonck [25]. Note that in an $n$-state DFA, the number of transitions is $O(n)$. We assume that the minimization is applied to each DFA or DFAO created without mentioning the minimization step every time.
@@
which is translated into an automaton. We know the automaton for $M_{\text{add}}$ has $2$ states and by Lemma 23, the automaton $M_{=(c/2)}$ has $\lfloor\log_{k}(c/2)\rfloor$ + 2 states. First we apply the product construction to $M_{=(c/2)}$ and $M_{\text{add}}$ such that the $M_{=(c/2)}$ part processes $y$ and the $M_{\text{add}}$ part processes $y\times y\times x$; the resulting automaton has $2(\lfloor\log_{k}(c/2)\rfloor+2)$ states. Each state in the product automaton is of the form $[q,p]$ where on input $x\times y$, $q$ is a state from the minimal automaton $M_{=(c/2)}$ representing the number of letters from $(c/2)_{k}$ matched with $y$ by Lemma 23 and $p\in\{0,1\}$ is a state from $M_{\text{add}}$ equal to $x]_{k}-2[y]_{k}$. Then we apply the $\exists y$ quantifier by removing the component corresponding to $y$ from the transitions and we get an NFA. If on input $x$ this NFA reaches $[q,p]$, from $q$ and $p$ we can determine how many letters from $x$ match $(c)_{k}$. So each state in this NFA corresponds to a state in the automaton in the proof of Lemma 23 for constant $c$. Then we do a subset construction to create a DFA from the NFA and the resulting DFA is the automaton from Lemma 23 with $O(\log c)$ states.
@@
First, we apply the $\land$ by product construction for $M_{\text{add}}$ and $M_{=c}$ such that the $M_{\text{add}}$ part processes $x\times y\times z$ and the $M_{=c}$ part processes $y$. The resulting automaton $M^{\prime}_{c}$ has $O(\log c)$ states. Let $\operatorname{Pref}_{k}(c,i)$ denote the first $i$ letters in $(c)_{k}$. On input $x\times y\times z$ where $y$ is $0^{*}y^{\prime}$ and $y^{\prime}$ is a prefix of $(c)_{k}$, the automaton $M^{\prime}_{c}$ leads to a state $[q,p]$ where $q\in\{0,1\}$ is from $M_{\text{add}}$ and $p=|y^{\prime}|$ is from $M_{=c}$ indicating how many letters from $y$ have been matched with $(c)_{k}$. Therefore, we have $q=[z]_{k}-([x]_{k}+[y]_{k})$ and $\operatorname{Pref}_{k}(c,p)]_{k}=[y]_{k}$. So $z]_{k}-[x]_{k}=q+[\operatorname{Pref}_{k}(c,p)]_{k}$ and to each state $[q,p]$ in $M^{\prime}_{c}$ a difference $z]_{k}-[x]_{k}$ can be attributed.
Then to implement the $\exists y$ quantifier, we remove the $y$ component from all transitions. We determinize the automaton to obtain $M_{c}$. Consider a (non-dead) state reachable on input $x\times z$ in $M_{c}$. This state consists of all $[q,p]$ such that $z]_{k}-[x]_{k}=q+[\operatorname{Pref}_{k}(c,p)]_{k}$. Since $q\in\{0,1\}$, there are at most two such $[q,p]$. Furthermore, each $[q,p]$ from $M^{\prime}_{c}$ appears in at most one state in $M_{c}$. Therefore, there are $O(\log c)$ states in $M_{c}$. The resulting automaton $M_{c}=(Q,\Sigma^{2}_{k},\delta,q_{0},A)$ can be formally defined as follows.
$Q$ $=\{d\in\mathbb{N}\ :\ d\in[0,c+1]\},$
@@
Note that not all $d\in[0,c+1]$ are reachable from the initial state.
@@
If $n$ is odd, we recursively compute a DFA $M_{n-1}(x,y)$ and use the expression
@@
In both cases, the translation could conceivably generate a nondeterministic automaton (by the projection). In this case, the subset construction is used to determinize the resulting automaton. Thus, at least in principle, this particular translation mechanism could take exponential time. However, we show that in this case, the subset construction does not cause an exponential blowup in the number of states.
@@
Let $n\geq 1$ be a fixed constant. The DFA $M_{n}$ with msd-first input for recognizing $y]_{k}=n[x]_{k}$ constructed recursively has $n$ states corresponding to differences $y]_{k}-n[x]_{k}$ in the range [0,n1][0,n-1].
@@
We prove the result by induction on $n$.
@@
We next remove the $y$ component from transitions to get the $n$-state automaton $M_{n}=(Q,\Sigma^{2}_{k},\delta,q_{0},A)$ as follows.
@@
Projecting away the $y$ component from transitions could have caused nondeterminism, but this is not the case. The differences represented by states are distinct and the output of transition function is no more than one state.
@@
First, we create an automaton $M^{\prime}_{n}$ from the usual product construction for $M_{n-1}$ and $M_{\text{add}}$. So this automaton has 2(n1)2(n-1) states. Consider an input $x\times y\times z$ for $M^{\prime}_{n}$ leading to some state $[q,p]$ where $q$ is from $M_{n-1}$ and $p$ is from $M_{\text{add}}$. By the construction of $M^{\prime}_{n}$, we know $q=[y]_{k}-(n-1)[x]_{k}$ and $p=[z]_{k}-([x]_{k}+[y]_{k})$. So we have $q+p=[z]_{k}-n[x]_{k}$. Therefore, to each state in the automaton $M^{\prime}_{n}$, a difference $z]_{k}-n[x]_{k}$ can be attributed. Furthermore, since $p\in\{0,1\}$ and $q\in\{0,\ldots,n-2\}$, all differences in $\{0,\ldots,n-1\}$ are present in $M^{\prime}_{n}$. Note that the difference 0 corresponds to exactly one pair—namely [0,0][0,0]—and the difference n1n-1 corresponds to only one pair—namely [n2,1][n-2,1]. All the other differences $d$ (those in $\{1,2,\ldots,n-2\}$) are represented by two states—namely, [d1,1][d-1,1] and $[d,0]$.
Next, we project away the $y$ component. The resulting automaton is nondeterministic, and we use the subset construction to create a deterministic automaton $M_{n}$. Consider a state from $M_{n}$ reachable on $x\times z$. This state consists of all pairs $[q,p]$ from $M^{\prime}_{n}$ such that $z]_{k}-n[x]_{k}=q+p$. As mentioned above, if $z]_{k}-n[x]_{k}$ is 0 or n1n-1, then there is only one such pair; otherwise, there are two such pairs. Furthermore, no $[q,p]$ appears in more than one state of $M_{n}$. So $M_{n}$ has $n$ states and the determinization did not cause a blow up in the number of states. The DFA $M_{n}=(Q,\Sigma^{2}_{k},\delta,q_{0},A)$ can be formally defined as follows.
@@
Theorem 27.
Proof.
Let $T(n)$ be the time used by the algorithm above to create $M_{n}$. We now analyze $T(n)$ based on the information from the proof of Theorem 26.
In Case 1, the number of transitions traversed to obtain $M_{n}$ is as follows: $4$ transitions for $M_{\text{add}}$, $nk^{3}$ transitions for the product construction of $M_{n/2}$ and $M_{\text{add}}$, $nk^{3}$ transitions for projecting away the $y$ component, $nk^{2}$ transitions for the determinized automaton, giving a total of $O(n)$ transitions in addition to the number of transitions traversed to create $M_{n/2}$.
In Case 2, the number of transitions traversed to obtain $M_{n}$ is as follows: $4$ transitions for $M_{\text{add}}$, $2(n-1)k^{3}$ transitions for the product construction of $M_{n-1}$ and $M_{\text{add}}$, $2(n-1)k^{3}$ transitions for projecting away the $y$ component, $nk^{2}$ transitions for the determinized automaton, giving a total of $O(n)$ transitions in addition to the number of transitions to create $M_{n-1}$.
@@
Let $n\geq 1$, $c\geq 0$ be fixed constants. For large enough values of $n$ and $c$, the automaton recognizing the relation $n[x]_{k}+c=[z]_{k}$ with msd-first input can be computed in $O(\log^{2}(c)\log\log(c)+n\log^{2}(n)+n\log(c)\log(n\log(c)))$ time.
@@
We first apply the product construction to $M_{n}$ and $M_{c}$. The resulting automaton $M^{\prime}_{n,c}$ has $O(n\log c)$ states.
Consider an input $x\times y\times z$ for $M^{\prime}_{n,c}$ leading to some state $[q,p]$ where $q$ is from $M_{n}$ and $p$ is from $M_{c}$. By the construction of $M^{\prime}_{n,c}$, we know that $q=[y]_{k}-n[x]_{k}$ and $p=[z]_{k}-[y]_{k}$. So we have $q+p=[z]_{k}-n[x]_{k}$. Therefore, to each state in the automaton $M^{\prime}_{n,c}$ a different $z]_{k}-n[x]_{k}$ can be attributed. Furthermore, since $q\in\{0,\ldots,n-1\}$ and $p\in[0,c+1]$, the attributed differences are in the range $[0,n+c]$. If we choose a difference $d\in[0,n+c]$, then we can exactly tell what combinations of $[q,p]$ correspond to $d$ (if any) and no $[q,p]$ corresponds to more than one $d$.
Next, we project away the $y$ component from $M^{\prime}_{n,c}$. The resulting NFA needs to be determinized by subset construction to get the DFA $M_{n,c}$. Consider a state from $M_{n,c}$ reachable on $x\times z$. This state consists of all $[q,p]$ from $M^{\prime}_{n,c}$ such that $q+p=[z]_{k}-n[x]_{k}$. As discussed above, the states from $M^{\prime}_{n,c}$ are assigned to only one state in $M_{n,c}$ and each state in $M_{n,c}$ is a difference $z]_{k}-n[x]_{k}$ in range $[0,n+c]$. The DFA $M_{n,c}=(Q,\Sigma^{2}_{k},\delta,q_{0},A)$ can be formally defined as follows.
$Q$ $=\{d\in\mathbb{N}\ :\ d\in[0,n+c]\},$
@@
Taking into account the minimization steps and the time spent on creating $M_{n}$ and $M_{c}$ based on the proof of Theorems 27 and 25, the time spent on creating $M_{n,c}$ is $O(\log^{2}(c)\log\log(c)+n\log^{2}(n)+n\log(c)\log(n\log(c))+(n+c)\log(n+c))$. So for large enough values of $n$ and $c$, the time spent on creating $M_{n,c}$ is $M_{n,c}$ is $O(\log^{2}(c)\log\log(c)+n\log^{2}(n)+n\log(c)\log(n\log(c)))$. ∎
@@
Let $n\geq 1$, $c\geq 0$ be fixed constants. For msd-first input, given an $m$-state DFAO generating $(h(i))_{i\geq 0}$, for large enough values of $n$ and $c$, we can compute the DFAO generating $(h(ni+c))_{i\geq 0}$ in $O(\log^{2}(c)\log\log(c)+n\log^{2}(n)+n\log(c)\log(n\log(c))+m^{2}(n+c)\log(m^{2}(n+c)))$ time.
@@
We first use the product construction for $M_{n,c}$ and $M_{h}$ such that the $M_{n,c}$ part processes the input $x\times y$ and the $M_{h}$ part processes the input $y$ to get the automaton $M^{\prime}$. Consider a state $[q,p]$ reachable on $x\times y$ in $M^{\prime}$ where $q$ is from $M_{n,c}$ and $p$ is from $M_{h}$. We have $q=[y]_{k}-n[x]_{k}\in[0,n+c]$ and $p=\delta_{h}(q_{0,h},y)$. This automaton has $m(n+c+1)$ states.
Next, we project the $y$ component away and use the subset construction to construct the DFAO $M$. On input $x$ the automaton reaches a state consisting of all pairs $[q,p]$ where $q\in[0,n+c]$ and $p=\delta_{h}(q_{0,h},(n[x]_{k}+q)_{k})$. Among these pairs there is one pair $q^{\prime},p^{\prime}]$ such that $q^{\prime}=c$ and we have $M(x)=\tau_{h}(p^{\prime})$. The DFAO $M=(Q,\Sigma_{k},\delta,q_{0},\Delta,\tau)$ can be formally defined as follows. Let $l=n+c$.
$Q$ $=\{[[0,p_{0}],\ldots,[l,p_{l}]]:p_{j}\in Q_{h}\},$
$\delta([[0,p^{\prime}_{0}],\ldots,[l,p^{\prime}_{l}]],a)$ $=[[0,p_{0}],\ldots,[c,p_{l}]]$
$\text{where}\ p_{j}=\delta_{h}(p^{\prime}_{j^{\prime}},b),j^{\prime}=\lfloor(na+j)/k\rfloor,b=(na+j)\bmod k,$
$q_{0}$ $=[[0,q_{0,h}],\ldots,[l,\delta_{h}(q_{0,h},(l)_{k})]],$
$\Delta$ $=\Delta_{h},$
$\tau([[0,p_{0}],\ldots,[l,p_{l}]])$ $=\{\tau_{h}(p_{j})\ :\ j=c\}.$
We need to know that the transition function is well-defined and has the correct output. The proof is similar to the proof of Theorem 10.
In each state of $M$, the first components of the pairs are consecutive differences in the range $[0,n+c]$ and the $p$ values are a subsequence of the interior sequence $(h^{\prime}(i))_{i\geq 0}$. Therefore, there are $O(m^{2}(n+c))$ states in $M$. Taking into account the minimization steps and the time spent on creating $M_{n,c}$ based on the proof of Theorem 28, for large enough values of $n$ and $c$, creating $M$ given $M_{h}$ takes $O(\log^{2}(c)\log\log(c)+n\log^{2}(n)+n\log(c)\log(n\log(c))+m^{2}(n+c)\log(m^{2}(n+c)))$ time. ∎
@@
Next, we project the $z$ component away and use the subset construction to construct the DFAO $M$. On input $x\times y$, the automaton $M$ reaches a state consisting of $0,p_{0}]$ and $1,p_{1}]$ where $p_{q}=\delta_{h}(q_{0,h},(q+[x]_{k}+[y]_{k})_{k})$ and we have $M(x,y)=\tau_{h}(p_{0})$.
The DFAO $M=(Q,\Sigma_{k},\delta,q_{0},\Delta,\tau)$ can be formally defined as follows.
$Q$ $=\{[[0,p_{0}],[1,p_{1}]]\ :\ p_{0},p_{1}\in Q_{h}\},$
[…]
$q_{0}$ $=[[0,q_{0,h}],[1,\delta_{h}(q_{0,h},1)]],$
$\Delta$ $=\Delta_{h},$
$\tau([[0,p_{0}],[1,p_{1}]])$ $=\tau_{h}(p_{0}).$
In each state of $M$, the $p_{0},p_{1}$ are a subsequence of the interior sequence $(h^{\prime}(i))_{i\geq 0}$. Therefore, there are $O(m^{2})$ reachable states in $M$. Taking minimization steps into account, creating $M$ given $M_{h}$ takes $O(m^{2}\log m)$ time. ∎

### 02 Apr 2026 v5

@@ 5 Construction by Büchi Arithmetic
Here the total number of transitions of an automaton $M$ as a measure for the computational complexity of constructing $M$. Some of the complexity bounds in this section include the addition of $O(\log i)$ or $O(i^{j})$ components for some constant $i$ which are not significant compared to multiplication in $O(\log i)$ or $O(i^{j})$.
@@
In Walnut each DFA or DFAO created is minimized. If $n$ is the number of states in an automaton (and the number of transitions is $O(n)$), the minimization steps in the implementation can use an algorithm running in $O(n\log n)$ time by Hopcroft [17] or Valmari [26]. The algorithm by Hopcroft can be applied to DFAOs with slight modifications as described by van Spaendonck [25]. Here we sometimes only consider the effects of a partial minimization to facilitate analysis in the following theorems; however, the upper bound on state complexity and runtime complexity is correct regardless and the runtime complexity computed takes into account the computational cost of a full minimization.
@@
which is translated into an automaton. We know the automaton for $M_{\text{add}}$ has $2$ states and by Lemma 22, the automaton $M_{=(c/2)}$ has $\lfloor\log_{k}(c/2)\rfloor$ + 2 states. First we apply the product construction to $M_{=(c/2)}$ and $M_{\text{add}}$ such that the $M_{=(c/2)}$ part processes $y$ and the $M_{\text{add}}$ part processes $y\times y\times x$; the resulting automaton has $2(\lfloor\log_{k}(c/2)\rfloor+2)$ states. Each state in the product automaton is of the form $[q,p]$ where on input $x\times y$, $q$ is a state from the minimal automaton $M_{=(c/2)}$ representing the number of letters from $(c/2)_{k}$ matched with $y$ by Lemma 22 and $p\in\{0,1\}$ is a state from $M_{\text{add}}$ equal to $x]_{k}-2[y]_{k}$. Then we apply the $\exists y$ quantifier by removing the component corresponding to $y$ from the transitions and we get an NFA. In the NFA there are no two same transitions from a single state leading to different states. So after subset construction the resulting DFA is the same as the NFA. Then we do minimziation and the resulting DFA is the automaton from Lemma 22 with $O(\log c)$ states.
@@
First, we apply the $\land$ by product construction for $M_{\text{add}}$ and $M_{=c}$ such that the $M_{\text{add}}$ part processes $x\times y\times z$ and the $M_{=c}$ part processes $y$. The resulting automaton $M^{\prime}_{c}$ has $O(\log c)$ states. Let $\operatorname{Pref}_{k}(c,i)$ denote the first $i$ letters in $(c)_{k}$. On input $x\times y\times z$ where $y$ is $0^{*}y^{\prime}$ and $y^{\prime}$ is a prefix of $(c)_{k}$, the automaton $M^{\prime}_{c}$ leads to a state $[q,p]$ where $q\in\{0,1\}$ is from $M_{\text{add}}$ and $p=|y^{\prime}|$ is from $M_{=c}$ indicating how many letters from $y$ have been matched with $(c)_{k}$. Therefore, we have $q=[z]_{k}-([x]_{k}+[y]_{k})$ and $\operatorname{Pref}_{k}(c,p)]_{k}=[y]_{k}$. So $z]_{k}-[x]_{k}=q+[\operatorname{Pref}_{k}(c,p)]_{k}$ and to each state $[q,p]$ reachable on $x\times y\times z$ in $M^{\prime}_{c}$ a difference $d=[z]_{k}-[x]_{k}$ can be attributed.
Then we apply the $\exists y$ quantifier by removing the component corresponding to $y$ from the transitions and we get an NFA. In the NFA there are no two same transitions from a single state leading to different states. So after subset construction the resulting DFA is the same as the NFA. Therefore, there are $O(\log c)$ states in the DFA and to each state of the DFA reachable on $x\times z$ a difference $d=[z]_{k}-[x]_{k}$ can be attributed.
From the proof of Theorem 3, we know the automaton recognizing $x]_{k}+c=[z]_{k}$ on input $x\times z$ only needs to keep track of a difference $z]_{k}-[x]_{k}$ in range $[0,c]$. So after some minimization, the resulting automaton $M_{c}=(Q,\Sigma^{2}_{k},\delta,q_{0},A)$ can be formally defined as follows:
$Q$ $=\{d\in\mathbb{N}\ :\ d\in[0,c]\},$
@@
Note that not all $d\in[0,c]$ are present in the minimized automaton.
@@
If $n$ is odd, we recursively compute a DFA $M_{n-1}(x,y)$ recognizing $y]_{k}=(n-1)[x]_{k}$ and use the expression
@@
In both cases, the translation can potentially generate a nondeterministic automaton (by the $\exists y$ projection). In this case, the subset construction is used to determinize the resulting automaton. Thus, at least in principle, this particular translation mechanism could take exponential time. However, we show that in this case, the subset construction does not cause an exponential blowup in the number of states.
@@
First by induction on $n$ we prove that the we can recursively create DFA $M_{n}$ with msd-first input recognizing $y]_{k}=n[x]_{k}$ with at most $n$ states.
@@
Then we apply the $\exists y$ quantifier by removing the component corresponding to $y$ from the transitions and we get an NFA. In the NFA there are no two same transitions from a single state leading to different states. So after subset construction the resulting DFA is the same as the NFA and we get the $n$-state automaton $M_{n}=(Q,\Sigma^{2}_{k},\delta,q_{0},A)$ as follows.
@@
Projecting away the $y$ component from transitions could have caused nondeterminism, but this is not the case.
@@
First, we create an automaton $M^{\prime}_{n}$ from the usual product construction for $M_{n-1}$ and $M_{\text{add}}$. So this automaton has 2(n1)2(n-1) states. Consider an input $x\times y\times z$ for $M^{\prime}_{n}$ leading to some state $[q,p]$ where $q$ is from $M_{n-1}$ and $p$ is from $M_{\text{add}}$. By the construction of $M^{\prime}_{n}$, we know $q=[y]_{k}-(n-1)[x]_{k}$ and $p=[z]_{k}-([x]_{k}+[y]_{k})$. So we have $q+p=[z]_{k}-n[x]_{k}$. Therefore, to each state in the automaton $M^{\prime}_{n}$ reachable on $x\times y\times z$, a difference $d=[z]_{k}-n[x]_{k}$ can be attributed. Furthermore, since $p\in\{0,1\}$ and $q\in\{0,\ldots,n-2\}$, all differences in $\{0,\ldots,n-1\}$ are present in $M^{\prime}_{n}$. Note that the difference 0 corresponds to exactly one pair—namely [0,0][0,0]—and the difference n1n-1 corresponds to only one pair—namely [n2,1][n-2,1]. All the other differences $d$ (those in $\{1,2,\ldots,n-2\}$) are represented by two states—namely, [d1,1][d-1,1] and $[d,0]$.
Next, we project away the $y$ component. The resulting automaton is an NFA, and we use the subset construction to create a DFA. Consider a state from $M_{n}$ reachable on $x\times z$. This state consists of some pairs $[q,p]$ from the NFA such that $z]_{k}-n[x]_{k}=q+p$. As mentioned above, if $z]_{k}-n[x]_{k}$ is 0 or n1n-1, then there is only one such pair; otherwise, there are two such pairs. So the DFA has $O(n)$ states and the determinization did not cause a blow up in the number of states.
From the proof of Theorem 6, we know the automaton recognizing $n[x]_{k}=[z]_{k}$ on input $x\times z$ only needs to keep track of a difference $z]_{k}-n[x]_{k}$ in range [0,n1][0,n-1]. So after some minimization, the DFA $M_{n}=(Q,\Sigma^{2}_{k},\delta,q_{0},A)$ can be formally defined as follows.
@@
Let $T(n)$ be the time used by the algorithm above to create $M_{n}$. We now analyze $T(n)$ based on the information above.
In Case 1, the number of transitions traversed for creating $M_{n}$ is as follows: $O(1)$ transitions for $M_{\text{add}}$, $O(n)$ transitions for the product construction of $M_{n/2}$ and $M_{\text{add}}$, $O(n)$ transitions for projecting away the $y$ component, $O(n)$ transitions for the determinized automaton, giving a total of $O(n)$ transitions in addition to the number of transitions traversed to create $M_{n/2}$.
In Case 2, the number of transitions traversed for creating $M_{n}$ is as follows: $O(1)$ transitions for $M_{\text{add}}$, $O(n)$ transitions for the product construction of $M_{n-1}$ and $M_{\text{add}}$, $O(n)$ transitions for projecting away the $y$ component, $O(n)$ transitions for the determinized automaton, giving a total of $O(n)$ transitions in addition to the number of transitions to create $M_{n-1}$.
@@
Let $n\geq 1$, $0\leq c<n$ be fixed constants. The automaton recognizing the relation $n[x]_{k}+c=[z]_{k}$ with msd-first input can be computed in $O(n\log(n)\log(n\log(n)))$ time.
@@
We first apply the product construction to $M_{n}$ and $M_{c}$. The resulting automaton $M^{\prime}_{n,c}$ has $O(n\log c)=O(n\log n)$ states.
Consider an input $x\times y\times z$ for $M^{\prime}_{n,c}$ leading to some state $[q,p]$ where $q$ is from $M_{n}$ and $p$ is from $M_{c}$. By the construction of $M^{\prime}_{n,c}$, we know that $q=[y]_{k}-n[x]_{k}$ and $p=[z]_{k}-[y]_{k}$. So we have $q+p=[z]_{k}-n[x]_{k}$. Therefore, to each state in the automaton $M^{\prime}_{n,c}$ reachable on $x\times y\times z$ a difference $d=[z]_{k}-n[x]_{k}$ can be attributed. Furthermore, since $q\in\{0,\ldots,n-1\}$ and $p\in\{0,\ldots,c\}$, the attributed differences are in the range [0,n+c1][0,n+c-1]. If we choose a difference $d\in[0,n+c-1]$, then we can exactly tell what combinations of $[q,p]$ correspond to $d$ (if any) and no $[q,p]$ corresponds to more than one $d$.
Consider all the states $0,d],[1,d-1],\ldots,[d,0]$ corresponding to the difference $d$, assuming all these states exist in $M^{\prime}_{n,c}$. If one of these states [q,dq][q,d-q] is reached on input $x\times y\times z$, then the state [q+i,dqi][q+i,d-q-i] is reached on input $x\times([y]_{k}+i)\times z$. Recall that we established we create automata so that they have the correct output regardless of the number of leading zeros read.
Next, we project away the $y$ component from $M^{\prime}_{n,c}$. The resulting NFA is determinized by subset construction to get a DFA. Consider a state from the DFA reachable on $x\times z$. This state consists of some $[q,p]$ from the NFA such that $q+p=[z]_{k}-n[x]_{k}$. Based on the discussion above, there are $d=O(n+c)$ states in the DFA. So the DFA has $O(n+c)=O(n)$ states.
From the proof of Theorem 6, we know the automaton recognizing $n[x]_{k}+c=[z]_{k}$ on input $x\times z$ for $0\leq c<n$ only needs to keep track of a difference $z]_{k}-n[x]_{k}$ in the range [0,n1][0,n-1]. So after some minimization, the DFA $M_{n,c}=(Q,\Sigma^{2}_{k},\delta,q_{0},A)$ can be formally defined as follows.
$Q$ $=\{d\in\mathbb{N}\ :\ d\in[0,n-1]\},$
@@
and $M_{n,c}$ has $O(n)$ states. Taking into account the minimization steps and the time spent on creating $M_{n}$ and $M_{c}$ based on the proof of Theorems 25 and 24 and considering $c=O(n)$, the time spent on creating $M_{n,c}$ is $O(n\log(n)\log(n\log(n)))$. ∎
@@
Let $n\geq 1$, $0\leq c<n$ be fixed constants. For msd-first input, given an $m$-state DFAO generating $(h(i))_{i\geq 0}$, we can compute the DFAO generating $(h(ni+c))_{i\geq 0}$ in $O(n\log n\log(n\log n)+nm^{2}\log(nm^{2}))$ time.
@@
We first use the product construction for $M_{n,c}$ and $M_{h}$ such that the $M_{n,c}$ part processes the input $x\times y$ and the $M_{h}$ part processes the input $y$ to get the automaton $M^{\prime}$. Consider a state $[q,p]$ reachable on $x\times y$ in $M^{\prime}$ where $q$ is from $M_{n,c}$ and $p$ is from $M_{h}$. We have $q=[y]_{k}-n[x]_{k}\in[0,n-1]$ and $p=\delta_{h}(q_{0,h},y)$. This automaton has $nm$ states.
Next, we project the $y$ component away and use subset construction to construct a DFAO. On input $x$ the automaton reaches a state consisting of some pairs $[q,p]$ where $q\in[0,n-1]$ and $p=\delta_{h}(q_{0,h},(n[x]_{k}+q)_{k})$. The $q$ are all values in the range [0,n1][0,n-1] and the $p$ values are length-$n$ subwords of $(h^{\prime}(i))_{i\geq 0}$. Furthermore, the output of the state is $\tau_{h}(p)$ where $p$ is from the state $[c,p]$ of $M^{\prime}$. So based on Theorem 1, there are $O(nm^{2})$ states in the automaton.
Taking into account the minimization steps and the time spent on creating $M_{n,c}$ based on the proof of Theorem 26, creating $M$ given $M_{h}$ takes $O(n\log n\log(n\log n)+nm^{2}\log(nm^{2}))$ time. ∎
@@
Next, we project the $z$ component away and use the subset construction to construct a DFAO. On input $x\times y$, the DFAO reaches a state consisting of some $[q,p]$ where $q\in\{0,1\}$ and $p=\delta_{h}(q_{0,h},(q+[x]_{k}+[y]_{k})_{k})$. The $p$ values are length-$2$ subwords of $(h^{\prime}(i))_{i\geq 0}$. So the DFAO has $O(m^{2})$ states.
Taking minimization steps into account, creating $M$ given $M_{h}$ takes $O(m^{2}\log m)$ time. ∎